# Fishers Harvest Parallel Unlearning in Inherited Model Networks

### v1

@@ Abstract.
Unlearning in various learning frameworks remains challenging, with the continuous growth and updates of models exhibiting complex inheritance relationships. This paper presents a novel unlearning framework, which enables fully parallel unlearning among models exhibiting inheritance. A key enabler is the new Unified Model Inheritance Graph (UMIG), which captures the inheritance using a Directed Acyclic Graph (DAG). Central to our framework is the new Fisher Inheritance Unlearning (FIUn) algorithm, which utilizes the Fisher Information Matrix (FIM) from initial unlearning models to pinpoint impacted parameters in inherited models. By employing FIM, the FIUn method breaks the sequential dependencies among the models, facilitating simultaneous unlearning and reducing computational overhead. We further design to merge disparate FIMs into a single matrix, synchronizing updates across inherited models. Experiments confirm the effectiveness of our unlearning framework. For single-class tasks, it achieves complete unlearning with 0% accuracy for unlearned labels while maintaining 94.53% accuracy for retained labels on average. For multi-class tasks, the accuracy is 1.07% for unlearned labels and 84.77% for retained labels on average. Our framework accelerates unlearning by 99% compared to alternative methods.
1. Introduction
The explosive growth of data volume and the increasing complexity of applications necessitate models that can constantly update to adapt to new data, constraints, and standards (Jayaraman and Evans, 2019). Accordingly, the need for data privacy, the auditing of illegal information from models, and compliance with regulations and industry standards arises. This leads to the concept of “model forgetting” or “unlearning,” which involves removing specific information from trained models.
The importance of unlearning lies in its ability to remove data that may raise privacy concerns, ensuring that the model can no longer recognize or rely on such data, in compliance with regulations such as the General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017). Additionally, unlearning plays a crucial role in data correction, removing the impact of erroneous or biased data from the model and enhancing its accuracy and fairness. This not only protects user privacy but also boosts the model’s quality and reliability, aligning with both ethical standards and legal requirements.
Machine unlearning becomes particularly challenging when models are consistently recorded with a directed acyclic graph (DAG) topology, where models inheriting from predecessors also require updates. Examples of such scenarios include a basic model being customized to meet various Machine Learning Operations (MLOps) lifecycle requirements (Matsui and Goya, 2023; Kreuzberger et al., 2023) or distilled to fit computational resource constraints (Polino et al., 2018).
The concept of model inheritance can be applied to prevalent learning frameworks, such as Federated Learning (FL) (Li et al., 2020a), Distributed Data-Parallel Learning (DDPL) (Li et al., 2020c), Incremental Learning (IL) (Van de Ven et al., 2022), and Transfer Learning (TL) (Weiss et al., 2016). In this case, unlearning even a single model would need to unlearn the entire subgraph rooted at the origin to ensure all descendant models are consistent and relevant.
@@ 1. Introduction
Research Gap. Various machine unlearning techniques have been proposed. Unfortunately, the inheritance relationships among related models prevent most existing unlearning methods from being executed in parallel, including re-training (Kim and Woo, 2022; Liu et al., 2024b), gradient ascent (Zhang et al., 2024), or knowledge distillation methods (Sinha et al., 2023). On the other hand, Sharding, Isolation, Slicing, and Aggregated Training (SISA) (Bourtoule et al., 2021) is a commonly used efficient method that reduces the dependence between data by dividing it into multiple segments and training multiple models in parallel. However, SISA is also inapplicable when there are inheritance relationships between models (Liu et al., 2024a). This is due to the fact that SISA needs to be performed sequentially from the starting point of updating to the end of the subgraph, incurring significant overhead and resource consumption during unlearning. In this sense, there is a pressing need for an effective and efficient way to perform unlearning in the presence of model inheritance relationships.
@@
Research Question. To address this research gap, we need to identify the commonalities across various prevalent frameworks in unlearning and determine how a new unlearning design can be universally applied. Specifically, we raise two key Research Questions (RQs):
@@
RQ1. How can unlearning requests be swiftly allocated to models that need updates and then prepare these models for unlearning tasks within a large-scale, interdependent model network?
@@
Contributions. This paper presents a novel parallel unlearning framework that can conduct efficient unlearning among large numbers of dependent models with complex inheritance relations, e.g., in FL, DDPL, IL, and TL. We first propose a new Unified Model Inheritance Graph (UMIG), to unveil the inheritance characteristic among models in various learning frameworks. With the assistance of the UMIG and Fisher Information Matrix (FIM), we develop the Fisher Inheritance Unlearning (FIUn) method that can efficiently pinpoint critical model parameters to be unlearned.
This allows independent unlearning adjustments on large-scale inherited models. To further enable effective one-shot unlearning of knowledge inherited from multiple upstream, we design a merging FIM function that consolidates disparate FIMs and identifies the critical parameters that need updating across all sources.
The contributions of this paper are summarized as follows:
  • We propose the UMIG, which abstracts various prevalent learning frameworks, e.g., FL, DDPL, IL, and TL, in a unified fashion using DAGs to represent models’ interdependence. This facilitates the analysis of large-scale inheritance-oriented model networks.
  • We design the FIUn method, which enables efficient parallel unlearning within the UMIG. The method leverages the FIM to quantify the significance of model parameters to unlearning tasks, and selectively scales the parameters according to their importance.
  • We further develop the merging FIM function to consolidate the FIMs from multiple upstream unlearning models into a cohesive matrix, thereby aggregating the unlearning tasks inherited from those models. This design facilitates the one-shot removal of inherited knowledge and reduces computational overhead within the UMIG.
Comprehensive experiments demonstrate that our method achieves complete unlearning for single-class unlearning tasks across all considered learning frameworks and model types, retaining an average accuracy of 94.53%. For multi-class unlearning tasks, the average unlearning accuracy drops to 1.07%, while the accuracy for retained labels reaches up to 84.77%. Additionally, our algorithm shows a 99% speed improvement across all considered frameworks compared to alternative methods.
Paper Organization. The rest of the paper is organized as follows. Section 2 reviews the preliminaries. Section 3 provides the proposed UMIG, followed by introducing the novel FIUn method and merging FIM function in Section 4. Section 5 presents the experimental settings. Section 6 presents our comprehensive experimental results based on four learning frameworks. Section 7 reviews the related works. Section 8 concludes this work.
2. Preliminaries
@@ 2.1. Machine Unlearning
Machine unlearning refers to removing the influence of specific data subsets from a trained model while maintaining its performance on the remaining data (Foster et al., 2024a; Umer and Polikar, 2021). In other words, the goal is to eliminate the impact of specific labels on the model’s predictions while ensuring that the prediction accuracy for the remaining labels and the overall performance of the model remain unaffected.
Let $D=\left\{\left(x_{i},y_{i}\right)\right\}^{N}_{i=1}$ represent a dataset where $x_{i}$ is the $i$-th training sample and $y_{i}\in\left\{1,\cdots,K\right\}$ is the corresponding class label. Let $C_{f}$ denote the set of labels to be unlearned, corresponding to the data set $D^{f}\subset D$ to be unlearned, and $C_{r}$ represent the set of retained labels, corresponding to the remaining data set $D^{r}=D\backslash D^{f}$.
Unlearning aims to eliminate the impact of $D^{f}$ with labels in $C_{f}$ from the model while preserving its performance on $D^{r}$ with labels in $C_{r}$. In this context, we define $\phi_{\theta}\left(\cdot\right):X\rightarrow Y$ as a function parameterized by $\theta\in\mathbb{R}^{m}$, where $X\in\mathbb{R}^{n}$ and $Y\in\mathbb{R}^{K}$. The $k$-th component of $\phi_{\theta}\left(x\right)$ represents the probability that sample $x$ belongs to class $k$. The objective is to adjust $\theta$ such that the model effectively “forgets” the influence of $C_{f}$ while retaining accurate predictions on $C_{r}$.
2.2. Fisher Information Matrix
The FIM is a useful tool in statistics for assessing the accuracy of parameter estimates (Kirkpatrick et al., 2017; Golatkar et al., 2020). It quantifies the impact of parameter changes on the probability distribution of observed data.
@@ 2.2. Fisher Information Matrix
Given a probability density function $p\left(x|\theta\right)$ conditioned on the model parameter $\theta$, an FIM $I\left(\theta\right)$ is the expected value of the second derivative of the negative log-likelihood function $\ell\left(\theta\right)$, as given by
@@
(1) $I\left(\theta\right)=\mathbb{E}\left[-\frac{\partial^{2}\ell\left(\theta\right)}{\partial\theta^{2}}\right],\ \text{where}\ \ell\left(\theta\right)=\ln p\left(x|\theta\right).$
@@
To simplify computation, in practice, the diagonal of the FIM can approximate the second derivative of the loss function (Steven, 1993; Aich, 2021). An equivalent form of $I\left(\theta\right)$ is given by
(2) $I(\theta)=\mathbb{E}\left[\left(\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)^{\top}\right)\right]\ .$
In this paper, the FIM is employed to quantify the importance of data to the model parameters. By comparing the FIMs of the training dataset and the to-be-unlearned dataset, we can identify the critical parameters related to the knowledge that needs to be unlearned.
3. Capturing Model-Inheritance: UMIG
@@
In this section, we explore the topological structure of four prevalent frameworks, FL, DDPL, IL, and TL, and analyze the impact of unlearning on the inherited models within these frameworks. We propose the UMIG, which can describe the model inheritance and update relationships in these learning frameworks in a unified manner.
@@
(c) IL and TL
(d) Decentralized FL
Figure 1. These diagrams demonstrate various learning frameworks that can be constituted to the proposed UMIG.
3.1. Diverse Learning Frameworks and Tasks
3.1.1. Federated Learning
A central server creates a global model, which participants download to locally update the model parameters based on their local data. The participants then send back their updated parameters. The central server aggregates the parameters to update and redistribute the global model. This process repeats until the model converges (McMahan et al., 2017; Li et al., 2020b; Krauß et al., 2024). As shown in Figure 1(a), the topology of FL consists of nodes representing model states and edges representing parameter transmission and aggregation. Updates to model $A$ trigger updates in the global model $D$ and then to $E$, $F$, $G$, and $H$. FL also has other structures.
@@ 3.1.1. Federated Learning
Multi-layer FL. This variant introduces an intermediate aggregation layer to traditional FL (Chu et al., 2024; Li et al., 2021), as depicted in Figure 1(b). This approach adds multiple servers that perform intermediate aggregation between the clients and the global server, thus reducing the burden on the global server, improving system scalability, and enhancing model training.
@@
DAG-FL. As shown in Figure 1(d), multiple clients update and aggregate models hierarchically in DAG-FL (Cao et al., 2021a, b; Yu et al., 2023). Each node can receive model updates from multiple parent nodes and send the updated results to multiple child nodes. This structure enables efficient model training and more flexible communication, better adapting to complex network environments and different application needs.
3.1.2. Distributed Data-Parallel Learning
@@
Training tasks are typically segmented into multiple sub-tasks, distributed among various computing nodes or servers for execution, as documented in (Ge et al., 2023; Dean et al., 2012; Abadi et al., 2016; Paszke et al., 2019; Li et al., 2014). Each node independently processes its designated sub-task and periodically synchronizes its model parameters with others to collectively refine the global model. This iterative process is crucial, particularly in DDPL topologies — similar to the FL topology depicted in Figure 1(a) — which require the sub-tasks $E$, $F$, and $H$, as well as the aggregated model $I$, to be updated continually with each new iteration of the global model $D$.
3.1.3. Incremental Learning
@@
Incremental (or “Continual”) learning entails training a model initially on a dataset and then progressively updating it as new data arrives, ensuring adaptability to the latest information (Han et al., 2023; Rajapaksha et al., 2023). IL can be visualized as shown in Figure 1(c), where nodes represent different updated states of the model, and edges depict the transition between model states after data updates. If the initial state model $G$ continues to update, then the subsequent models $A$ and $B$ that receive new data, as well as their related models, also need updates.
3.1.4. Transfer Learning
TL involves pre-training a base model on one dataset and then applying this model to a new task. Typically, this process includes fine-tuning the parameters of the last few layers to better suit the new task (Sendner et al., 2023; Rajapaksha et al., 2023). The model continues to train on the specific dataset of the new task to adapt fully to its requirements. Through ongoing fine-tuning and training, the model is optimized for the best performance on the new task. The TL topology, similar to the IL topology shown in Figure 1(c), also follows a comparable model updating process.
In these diverse learning frameworks, the inheritance relationship between models is pivotal. When a base model is updated or optimized, all corresponding child models derived from it must also be updated. To accurately capture this inheritance relationship, we employ the DAG topology as an effective abstraction method to capture the dependencies among models in these learning frameworks. The use of the DAG ensures that when a single model is updated, the entire subgraph—originating from this model and encompassing all related child models—is appropriately updated.
3.2. Unified Model Inheritance Graph (UMIG)

### v4

# Parallel Unlearning in Inherited Model Networks

@@ Abstract
Unlearning is challenging in generic learning frameworks with the continuous growth and updates of models exhibiting complex inheritance relationships. This paper presents a novel unlearning framework that enables fully parallel unlearning among models exhibiting inheritance. We use a chronologically Directed Acyclic Graph (DAG) to capture various unlearning scenarios occurring in model inheritance networks. Central to our framework is the Fisher Inheritance Unlearning (FIUn) method, designed to enable efficient parallel unlearning within the DAG. FIUn utilizes the Fisher Information Matrix (FIM) to assess the significance of model parameters for unlearning tasks and adjusts them accordingly. To handle multiple unlearning requests simultaneously, we propose the Merging-FIM (MFIM) function, which consolidates FIMs from multiple upstream models into a unified matrix. This design supports all unlearning scenarios captured by the DAG, enabling one-shot removal of inherited knowledge while significantly reducing computational overhead. Experiments confirm the effectiveness of our unlearning framework. For single-class tasks, it achieves complete unlearning with 0% accuracy for unlearned labels while maintaining 94.53% accuracy for retained labels. For multi-class tasks, the accuracy is 1.07% for unlearned labels and 84.77% for retained labels. Our framework accelerates unlearning by 99% compared to alternative methods.
I Introduction
In the rapidly evolving landscape of data and application complexity, models require frequent updates to adapt to new data, constraints, and standards, highlighting the critical need for robust data privacy measures, auditing of illicit information, and adherence to regulatory and industry standards [jayaraman]. The process of “model forgetting” or “unlearning” emerges as essential, involving the removal of specific information from trained models to address privacy concerns, ensuring models do not retain or utilize sensitive data in violation of regulations like the General Data Protection Regulation (GDPR) [voigt2017eu]. Moreover, unlearning facilitates data correction by eliminating the influence of erroneous or biased inputs, thereby not only protecting user privacy but also enhancing the model’s accuracy, fairness, and overall quality.
Many existing unlearning advances (e.g., [wang2025gru, yang2025exploring]) typically focus on removing the influence of specific data from a single model, without considering relationships among models. However, in an Inherited Model Network (IMN), where models are organized as a Directed Acyclic Graph (DAG), unlearning becomes significantly more complex: updates to one model must be consistently propagated to all its downstream inherited models. This structure introduces cascading effects and computational challenges that are unique to IMNs. This challenge arises in scenarios such as customizing a base model to meet various Machine Learning Operations (MLOps) lifecycle requirements [mbma] or distilling it to fit computational resource constraints [polino2018model]. These situations can be abstracted into prevalent learning frameworks such as Federated Learning (FL) [li2020review], Distributed Data-Parallel Learning (DDPL) [ddap], Incremental Learning (IL) [IL], and Transfer Learning (TL) [TL]. In such cases, unlearning even a single model necessitates unlearning the entire subgraph rooted at the origin to ensure that all descendant models remain consistent and relevant.
@@ I Introduction
Research gaps. Various machine unlearning techniques have been proposed. Unfortunately, the inheritance relationships among related models prevent most existing unlearning methods from being executed in parallel, including re-training [retrain], gradient ascent [ga], or knowledge distillation methods [distil1Unl]. On the other hand, Sharding, Isolation, Slicing, and Aggregated Training (SISA) [sisa, yu2024splitunlearning] is a commonly used efficient method that reduces the dependence between data by dividing it into multiple segments and training multiple models in parallel. However, SISA is also inapplicable when there are inheritance relationships between models [liu2024]. This is because SISA needs to be performed sequentially from the starting point of updating to the end of the subgraph, incurring significant overhead and resource consumption during unlearning. In this sense, there is a pressing need for an efficient (i.e., unlearning speed) and effective (i.e., accuracy) way to perform unlearning in the presence of model inheritance relationships.
@@
Research questions. To address this research gap, we need to identify the commonalities across various prevalent frameworks in unlearning and determine how a new unlearning design can be universally applied. We raise two key Research Questions (RQs):
@@
RQ1. How can unlearning requests be quickly allocated to models that need updates and what unique features could be realized during unlearning in a large-scale and interdependent model network?
@@
Contributions. This paper introduces a novel parallel unlearning framework for efficiently handling unlearning across numerous dependent models with complex inheritance relations, such as in FL, DDPL, IL, and TL. The contributions are summarized as follows:
  • $\bullet$
    We capture various unlearning scenarios in IMNs, which are chronologically DAG-based. By analyzing scenarios with single- and multi-root unlearning requests, and the impact of path depths and multi-path structures on unlearning, we provide precise and efficient identification of models to be unlearned.
  • $\bullet$
    We introduce the Fisher Inheritance Unlearning (FIUn) method to enable efficient parallel unlearning within DAGs. FIUn compares the Fisher Information Matrices (FIMs) of the initial unlearning models (upon the data to be unlearned) with the FIMs of all inherited models, identifies the model parameters that require updates, and then proportionally adjusts parameters in parallel across the inherited models based on their unlearning impact.
  • $\bullet$
    To handle multiple unlearning requests simultaneously, FIUn incorporates a Merging-FIM (MFIM) function that consolidates FIMs from multiple upstream models into a single unified matrix. This allows FIUn to address all unlearning scenarios, facilitating one-shot removal of inherited knowledge while greatly reducing computational overhead.
Extensive experiments indicates that our method significantly outperforms all comparison methods in terms of unlearning speed on widely-accepted public image and text datasets. Compared to re-training, our method achieves comparable levels of unlearning accuracy. Even on large-scale models, our method demonstrates a certain degree of effectiveness and applicability. A preliminary version of this work has been posted on arXiv. [^1] This work is based on a preprint post: https://arxiv.org/abs/2408.08493.
The rest of this paper is organized as follows. The related works are reviewed in Section II. The preliminaries are in Section III. The explored topological structure frameworks are discussed in Section IV. The designed parallel unlearning method in Section V. The experimental settings are presented in Section VI, followed by the evaluation in Section VII. Section VIII concludes this work.
II Related Work
Although FL [li2020federated], DDPL [ddap], IL [IL], and TL [TL] have addressed challenges such as data privacy [li2021survey], training efficiency [dean2012large], knowledge retention and unlearning [han2023anomaly], and cross-domain knowledge transfer [rajapaksha2023improving], systematic exploration combining these frameworks is limited. Most studies have only attempted simple method combinations without addressing their shared foundation model inheritance. This aspect is crucial in large-scale unlearning, where computational overhead and time consumption increase with inheritance depth, even when unlearning is parallelized at the same depth. This remains true across existing unlearning methods, which are discussed below.
Federated Unlearning. Wang et al. [wang2024server] proposed a server-initiated federated unlearning framework, SIFU, which identifies and quantifies the impact of low-quality client data. By combining gradient ascent with high-quality data augmentation, the method efficiently eliminates the negative influence of such data on the global model. Tao et al. [tao2024communication] introduced the Total Variation (TV) Stability to measure model sensitivity to small data perturbations and modified the FedAvg algorithm to improve communication efficiency while enabling precise client- and sample-level unlearning. To address performance degradation caused by gradient conflicts, Pan et al. [pan2025federated] proposed the FedOSD method, which designs a dedicated unlearning loss function and optimized the model along the steepest descent direction closest to the target client’s gradient, without interfering with others, achieving a balance between unlearning and utility. Gao et al. [gao2024verifi] introduced the VeriFi framework, a framework that supports active verification of unlearning effects and includes a more efficient unlearning algorithm, uS2U, along with two non-invasive verification methods, vFM and vEM, enhancing the security and verifiability of federated unlearning.
Distributed Data-Parallel Unlearning. In DDPL knowledge unlearning, handling the unlearning process is restricted to the data shard on the specific device, followed by model aggregation and iterative training. If unlearning is required from a specific shard, only the model trained on that shard is re-trained. Approximate machine unlearning directly modifies trained model parameters, employing methods such as 1) Gradient-based approaches [golatkar2020eternal], which update model gradients to gradually diminish the data’s influence, and 2) Data perturbation [tarun2023fast], which introduces noise to reduce specific data points’ impact on the model. These methods facilitate effective knowledge unlearning in DDPL.
Incremental Unlearning. Research on unlearning in IL is still a developing area, but some recent methods show promise. Projected-Gradient Unlearning (PGU) [hoang2024learn] calculates gradients for the unlearning dataset, projects them onto the Core Gradient Space (CGS) derived from the training set, and subtracts the projected component. The residuals, now orthogonal to the CGS, are used to update model parameters. Effective for one-time and batch-wise unlearning, PGU is well-suited for incremental learning processes. Zuo et al. [zuo2024ecil] proposed a class-level unlearning method called eCIL-MU, which uses embedding to map data into vectors and store them in a vector database. This enables efficient and non-destructive unlearning by reusing the structure of class-incremental learning tasks.
Transfer Unlearning. Research on unlearning specific target task knowledge has been relatively scarce in TL. To address this challenge, Sepahvand et al. [sepahvand2024data] proposed fine-tuning pre-trained models through data selection. Subsequently, the entire network was fine-tuned on the target task using gradient descent.
Machine Unlearning. Nguyen et al. [nguyen2020variational] proposed a loss function minimizing Kullback-Leibler divergence to ensure certified data removal in Bayesian models, using two techniques to correct posterior inaccuracies and prevent catastrophic forgetting for small data subsets. Ginart et al. [ginart2019making] introduced Q-k-means and DC-k-means algorithms for k-means clustering, using quantized cluster centers and divide-and-conquer to efficiently update models without retraining, ensuring precise data removal while maintaining clustering quality. Kurmanji et al. [kurmanji2023towards] presented SCRUB, a teacher-student model optimizing student weights to diverge from the teacher’s output on the forget dataset ($D_{f}$) while aligning with the retain dataset ($D_{r}$), balancing forgetting and utility for bias removal, confusion resolution, and privacy. Zhu et al. [zhu2024decoupling] proposed TARF, using annealed gradient ascent and target-aware gradient descent to decouple specific concepts, preserving performance on unaffected data. Both modified model weights for efficient approximate unlearning without strict certified removal.
Ours. Existing unlearning methods struggle with high computational costs, resource demands, and implementation challenges, particularly in large-scale settings. Our FIUn method enables fully parallel unlearning by adapting single-model techniques [foster2024fast, golatkar2020eternal] for IMNs, improving efficiency, resource utilization, and adaptability while achieving effectiveness comparable to re-training.
III Preliminaries
@@ III-A Machine Unlearning
Machine unlearning refers to removing the influence of specific data subsets from a trained model while maintaining its performance on the remaining data [foster2024zero].
Let $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ denote the full training dataset, where $x_{i}$ is the $i$-th input and $y_{i}\in\{1,\dots,K\}$ is its corresponding label. Let $D^{f}\subset D$ be the subset of samples to be unlearned (forget set), and $D^{r}=D\setminus D^{f}$ be the remaining samples (retain set). Let $\phi_{\theta}:X\rightarrow Y$ be the model parameterized by $\theta\in\mathbb{R}^{m}$, and $\phi_{\theta^{\prime}}$ be the updated model after unlearning. The unlearning process is considered successful if: (i) $\phi_{\theta^{\prime}}\perp D^{f}$, meaning that the predictions of $\phi_{\theta^{\prime}}$ are statistically independent of the removed data $D^{f}$, ideally resembling those of a model trained only on $D^{r}$; (ii) $L(\phi_{\theta^{\prime}},D^{r})\approx L(\phi_{\theta},D^{r})$, ensuring the model retains performance on the remaining data.
In practice, this entails adjusting the model weights $\phi_{\theta}$ to obtain $\phi_{\theta^{\prime}}$ that removes the influence of $D^{f}$ while preserving knowledge from $D^{r}$. Crucially, model parameters are inherently heterogeneous—different weights contribute unequally to different classes, samples, or tasks [koh2017understanding]. Effective unlearning therefore requires selective and targeted parameter adaptation. This motivates a range of gradient-based and sensitivity-aware approaches—including influence functions, Fisher Information, and other parameter importance metrics—that identify and update only the most relevant components in a principled and efficient manner.
III-B Fisher Information Matrix
@@
FIM is a useful tool in statistics for assessing the accuracy of parameter estimates [golatkar2020eternal]. It quantifies the impact of parameter changes on the probability distribution of observed data. Given a probability density function $p\left(x|\theta\right)$ conditioned on the model parameter $\theta$, an FIM $I\left(\theta\right)$ is the expected value of the second derivative of the negative log-likelihood function $\ell\left(\theta\right)$, as given by
@@
$I\left(\theta\right)=\mathbb{E}\left[-\frac{\partial^{2}\ell\left(\theta\right)}{\partial\theta^{2}}\right],\ \text{where}\ \ell\left(\theta\right)=\ln p\left(x|\theta\right).$ (1)
@@
To simplify computation, in practice, the diagonal of FIM can approximate the second derivative of the loss function [steven1993fundamentals]. An equivalent form of $I\left(\theta\right)$ is given by
$I(\theta)=\mathbb{E}\left[\left(\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)\left(\frac{\partial\ell\left(\theta\right)}{\partial\theta}\right)^{\top}\right)\right]\ .$ (2)
The FIM holds significant importance in the theory of machine unlearning. The natural gradient regards the FIM as a Riemannian metric tensor of the parameter space, used to define the optimal direction of parameter perturbation [amari1998natural]. Furthermore, the FIM is the expected form of the Hessian, aligning with the concept of influence functions for analyzing the impact of individual samples [koh2017understanding]. From an information-theoretic perspective, a larger FIM value indicates that the corresponding parameter carries more discriminative information about the training data and is therefore more critical to the model’s behavior.
An intuitive baseline is to simply remove the columns in the last FC layer weights corresponding to the classes to be unlearned. However, this approach has significant limitations: (i) it is only applicable to class-level unlearning and cannot be extended to sample-level unlearning tasks; (ii) simple eliminating weight columns disrupts the decision boundaries of the retained data, leading to substantial accuracy degradation [lee2023undo]; and (iii) in IMNs, this method only affects the current node and fails to ensure consistent propagation of unlearning information across all inherited models. To address these issues, this paper employs the FIM to quantify the importance of model parameters to the training data. By comparing the FIMs of the training dataset and the unlearning dataset, we can identify the critical parameters related to the knowledge to be unlearned.
IV Capturing Model-Inheritance
@@
In this section, we explore the topological structure of four prevalent frameworks, i.e., FL, DDPL, IL, and TL, and analyze the impact of unlearning on the inherited models within these frameworks. We propose using DAG, which can visualize and capture the model inheritance and update relationships in these learning frameworks in a unified manner.
IV-A Diverse Learning Frameworks and Tasks
Federated Learning. A central server creates a global model, which participants download to locally update the model parameters based on their local data. The participants then send back their updated parameters. The central server aggregates the parameters to update and redistribute the global model. This process repeats until the model converges [krauss2024automatic]. As shown in Fig. 1(a), the topology of FL consists of nodes representing model states and edges representing parameter transmission and aggregation. Updates to model $A$ trigger updates in the global model $D$ and then to $E$, $F$, $H$, and $I$.
@@ IV-A Diverse Learning Frameworks and Tasks
(c) Decentralized FL
(d) IL and TL
Figure 1: These diagrams show learning frameworks modeled as a DAG, where one user may create multiple models, and multiple models may be created by different users.
@@
Multi-Layer FL. This variant introduces an intermediate aggregation layer to traditional FL [li2021survey], as depicted in Fig. 1(b). This approach adds multiple servers that perform intermediate aggregation between the clients and the global server, thus reducing the burden on the global server, improving system scalability, and enhancing model training.
@@
Decenralized FL. As shown in Fig. 1(c), multiple clients update and aggregate models hierarchically in DAG-FL [yu2023ironforge]. Each node can receive model updates from multiple parent nodes and send the updated results to multiple child nodes. This structure enables efficient model training and more flexible communication, better adapting to complex network environments and different application needs.
@@
Distributed Data-Parallel Learning. Training tasks are divided into sub-tasks and distributed across computing nodes or servers [liu2024fishers, dean2012large]. Each node processes its sub-task independently and periodically synchronizes parameters to update the global model. This iterative process is essential in DDPL topologies—similar to the FL topology in Fig. 1(a)—where sub-tasks $E$, $F$, and $H$ and the aggregated model $I$ must be updated with every new iteration of the global model $D$.
@@
Incremental Learning. IL entails training a model initially on a dataset and then progressively updating it as new data arrives, ensuring adaptability to the latest information [han2023anomaly]. IL can be visualized as shown in Fig. 1(d), where nodes represent different updated states of the model, and edges depict the transition between model states after data updates. If the initial state model $G$ continues to update, then the subsequent models $A$ and $B$ that receive new data, as well as their related models, also need updates.
Transfer Learning. TL pre-trains a base model on one dataset and adapts it to a new task by fine-tuning, typically on the last layers, using task-specific data [rajapaksha2023improving]. Through continued fine-tuning, the model aligns with the new task and achieves optimal performance. The TL topology mirrors the IL topology in Fig. 1(d), following a similar update process.
In these learning frameworks, inheritance between models is crucial: when a base model is updated, all its child models must also be updated. We use a DAG topology to represent these dependencies, ensuring that updates to one model propagate through its entire subgraph of descendants.
IV-B Inherited Model Networks